# IMAIA: Interactive Maps AI Assistant for Travel Planning and Geo-Spatial Intelligence

### v3

@@ Abstract
Map applications are still largely point-and-click, making it difficult to ask map-centric questions or connect what a camera sees to the surrounding geospatial context with view-conditioned. We introduce IMAIA, an interactive Maps AI Assistant that enables natural-language interaction with both vector (street) maps and satellite imagery, and augments camera inputs with geospatial intelligence to help users understand the world. IMAIA comprises two complementary components. Maps Plus treats the map as first-class context by parsing tiled vector/satellite views into a grid-aligned representation that a language model can query to resolve deictic references (e.g., “the flower-shaped building next to the park in the top-right”). Places AI Smart Assistant (PAISA) performs camera-aware place understanding by fusing image–place embeddings with geospatial signals (location, heading, proximity) to ground a scene, surface salient attributes, and generate concise explanations. A lightweight multi-agent design keeps latency low and exposes interpretable intermediate decisions. Across map-centric QA and camera-to-place grounding tasks, IMAIA improves accuracy and responsiveness over strong baselines while remaining practical for user-facing deployments. By unifying language, maps, and geospatial cues, IMAIA moves beyond scripted tools toward conversational mapping that is both spatially grounded and broadly usable.
I Introduction
@@
Modern map applications remain largely point–and–click: users pan and zoom, then issue basic, limited and inflexible queries. This interaction model breaks down for map-centric, view-conditioned queries—e.g., “What’s the flower-shaped building next to the park in the top-right of what I’m viewing?”—and for connecting what a mobile camera sees to the surrounding geospatial context. The desire to explore and navigate unfamiliar environments is fundamental, yet current tools remain inadequate: traditional travel and mapping systems, constrained by static methodologies, struggle with real-world dynamism—fluctuating conditions, imprecise signals, and unexpected disruptions that degrade the experience. Meanwhile, travel planning, navigation, and local discovery are typically engineered as isolated modules, yielding fragmented interactions and brittle handoffs [1, 2]; prior work even documents user behavior under disrupted plans [3], yet practical systems still fail to support coherent, conversational map understanding.
@@
The recent rapid advancements in Large Language Models (LLMs) present a transformative opportunity to transcend these limitations [4, 5, 6]. LLMs possess an unprecedented ability to process and synthesize diverse multimodal inputs, including text, imagery, geospatial data, and contextual cues [7]. This capability is paving the way for a new generation of intelligent systems that are not only cohesive but also inherently adaptive [8]. This work is situated at the confluence of two pivotal trends in geospatial AI. Firstly, LLMs are increasingly adept at interpreting unstructured or ambiguous geospatial information, transforming vague user requests into precise map coordinates or deriving actionable insights from noisy datasets [8]. Secondly, advancements of conversational AI underscores the necessity of multi-turn, context-aware interactions for complex tasks such as navigation and discovery, where user needs and intentions can shift dynamically [9]. Beyond this, Vision–language models (VLMs [10]) can describe images, but without explicit spatial grounding to the current map state (viewport, scale, nearby entities) and geospatial signals (location, heading, proximity), responses are often brittle or slow. Rather than fragmenting planning, navigation, and local discovery into separate and loosely coupled modules, our approach unifies these capabilities within a single cohesive framework.
In this paper, we present the Interactive Maps AI Assistant (IMAIA), depicted in Figure 1. IMAIA is built around two tightly integrated components—Maps Plus and the Places AI Smart Assistantvthat operate in concert under a multi-agent orchestration layer. By coupling Maps Plus with PAISA, the system eliminates cross-module discontinuities and enables seamless reasoning across both map-centric spatial analysis and camera-based place understanding. This design supports fluid transitions between high-level reasoning on maps and fine-grained perception of real-world scenes, resulting in a coherent end-to-end user experience. The overarching goal of IMAIA is to deliver an interactive geospatial platform that bridges online trip planning with on-site exploration and discovery. By leveraging the complementary strengths of large language models (LLMs), vision–language models (VLMs), and a flexible multi-agent architecture, IMAIA provides a mapping experience that is not only more natural and adaptive but also capable of intelligently supporting both digital exploration and real-world navigation.
Maps Plus (map-centric grounding). We parse vector/satellite maps into a grid-aligned representation and align detected entities (e.g., roads, parks, water bodies) to a geospatial index, enabling efficient resolution of deictic, view-conditioned queries (e.g., “the lake in the top-right tile”) that support intelligent trip planning.
PAISA (multimodal understanding). PAISA fuses visual input from the camera with geospatial signals—including location, heading, and proximity—within a coordinated multi-agent architecture. By combining these heterogeneous modalities across orchestration, location intelligence, navigation, and spatial reasoning agents, the system delivers guidance that is both concise and context-aware, grounding real-world assistance in a rich multimodal understanding of the user’s environment.
IMAIA as a unified system. By tightly integrating Maps Plus and PAISA within a coordinated framework, IMAIA establishes an end-to-end pipeline that bridges map-based exploration with environment-grounded place understanding. Empirically, IMAIA achieves substantial improvements over strong baselines. Maps Plus raises place detection accuracy from under 43% to nearly 90%, while PAISA enables human-centered navigation that reduces detours compared to rigid turn-by-turn instructions. For spatial reasoning, our distilled model reaches 84% accuracy—a 3× gain over large multimodal LLMs—and delivers a 7.3 inference speedup (1.7s vs. 12.4s) over agent-based pipelines, ensuring responsiveness in real-world settings.
II Related Work
II-A Geospatial/Maps Intelligence with LLMs
@@
II-B Spatial Intelligence and Reasoning with VLMs
@@
III Interactive Maps AI Assistant
III-A Maps Plus
Travelers usually begin by asking broad questions such as “What are the top locations to visit in X?”, “What should I see in X?”, or “Plan me a route through X.” Previous studies have shown that a large language model (LLM) can answer such queries either directly from its internal memory or by coordinating with external tools [8, 5], as illustrated in Figure 3. These approaches are referred to as Single Model, Model + Location, and Model + Verbose Location. Once users begin examining an interactive map, however, they often pose richer, map-centric questions that require spatial reasoning about what they are actually viewing rather than a static list of attractions. To meet these needs, we introduce a multimodal system that blends an LLM with image input and geospatial search (User Interface as shown in Figure 2). A user can click on map tiles or satellite imagery and converse naturally about that view, receiving answers grounded in both textual knowledge and visual context. As shown in Figure 3, our proposed method first determines the geographic focus and zoom level of the user’s current view, then scans the surrounding imagery on a regular grid to extract visual features and detect salient geographic entities. Finally, it queries a geospatial index with those entities and synthesizes the results so the LLM can craft an informed, location-aware response. By treating the map itself as conversational context, the system supports fluid trip planning and exploratory tasks that go well beyond what text-only approaches can deliver.
III-A1 Location Awareness
@@
III-A2 Grid-Based Spatial Analysis
Next, a simplified map with a grid overlay is provided to multimodal LLM as shown in Figure 4. The model is tasked with identifying grid cells that contain significant map entities, such as roads, parks, or water bodies. This step allows for spatial correlation analysis by associating detected entities with their positions on the map. By segmenting the map in this way, the model can break down complex spatial relationships and make them more accessible for downstream tasks, such as answering questions about specific regions.
III-A3 Entity Search and Query Resolution
@@
III-B Places AI Smart Assistant (PAISA)
@@
III-B1 Interactive Navigation Agent
@@
III-B2 Location Intelligence Agent
@@
III-B3 Spatial Understanding Agent
Existing vision–language models fine-tuned for spatial reasoning (e.g., ASMv2 [17], SpatialVLM [18], SpatialRGPT [19]) often suffer from high latency and misalignment with the agent’s task requirements. To overcome these limitations, we distilled a Florence-2 model [21] from GPT-4o [22] with instruction fine-tuning, resulting in a more efficient and task-adapted spatial reasoning module. To further enhance the assistant’s spatial understanding, we introduce a spatial scene-graph understanding agent powered by a fine-tuned lightweight multimodal language model. This agent processes a single photo and identifies the most visually salient objects in the scene—those that are most recognizable or attention-catching (e.g., store signs, flags, building façades)—along with their spatial relationships, expressed either as a scene graph or in natural language. The extracted spatial information can be utilized in two complementary ways: (1) When the assistant retrieves cached street-level images of a destination, the agent generates spatially grounded descriptions that highlight the relative positions of salient objects, helping users more reliably recognize the destination in the real world. (2) When users upload a photo of their current surroundings, the agent analyzes the scene and produces a structured spatial representation, which then serves as contextual input for downstream agents. By grounding both destination imagery and user-provided photos in spatial relationships, the system offers guidance that is more interpretable, robust, and actionable.
The distillation pipeline comprises three stages: (i) entity mining, (ii) object and spatial-relation extraction, and (iii) supervised fine-tuning. In stage (i), we prompt GPT-4o-mini [23] to simulate an urban wayfinding scenario and, given a random street-view image, propose candidate “key” items. Repeating this five times per image over 40,000 street-view images, we retain the top-10 most frequently mentioned items per image as the “key” entities. In stage (ii), we localize each entity’s 2D position using YOLO-World [24] and estimate its 3D depth with Depth Anything V2 [25]. We then assemble a structured record of each entity’s 2D coordinates and distance to the camera, and pair the image with bounding-box overlays and a metric depth map to prompt the GPT-4o for pairwise spatial relationships. Prompts follow the Set-of-Mark Prompting paradigm to enhance visual grounding and reasoning. In stage (iii), we build a supervised fine-tuning dataset by pairing each annotated image with over 15 variants of spatial queries that reflect realistic urban wayfinding needs, such as relative positioning (“What object is directly to the left of the store sign?”) and orientation (“Which structure faces the street?”). Each query is matched with answers derived from the outputs of stage (ii), yielding diverse training samples that encourage the model to generalize beyond simple object detection. We fine-tune a Florence-2-large with full-parameter training, casting the task into its native Dense Region Captioning format so the model learns to generate spatially grounded, linguistically coherent descriptions conditioned on visual regions and relational context. This final stage enables the distilled model to inherit GPT-4o’s reasoning ability while remaining efficient and task-adapted, ultimately producing a lightweight yet robust spatial reasoning module capable of supporting real-time navigation.
III-B4 Handle complex query with multi-agent reasoning
@@
IV Results
IV-A Place detection accuracy with Maps Plus
In this section, we evaluate and compare the quality of geospatial entity search across several methods. Specifically, we benchmark maps plus approach against three widely used baselines (Figure 3) in Table I: (1) Single Model: The model receives only the user query and the current map view image, relying solely on its pretrained knowledge. (2) Model + Location: The input includes the query, the map view image, and the geographic coordinates (latitude and longitude) of the map view. (3) Model with Verbose Location: Similar to the previous setup, but the location input is replaced with verbose descriptors such as city names and landmarks. In this study, we construct a dataset by selecting 10 cities across the United States and sampling points of interest (POIs) within a 20-kilometer radius of each city center. For each POI, we employ GPT-4o to generate synthetic queries based on information such as the POI’s attributes, geographic coordinates, and related contextual data. This process yields a total of 4,300 queries, for example, What is the lake at the top left part of the map.
@@ IV-A Place detection accuracy with Maps Plus
Table I presents the accuracy achieved by various methods. The results indicate that our proposed approach attains significantly higher accuracy compared to the baselines, without requiring any fine-tuning of the LLM. The superior performance of the proposed method can be attributed to its efficient integration of grounding data, specifically entities retrieved from geospatial services, into the LLM’s processing. For instance, consider the query ”What is the coffee shop below the cinema?”. The proposed method not only furnishes a pertinent set of local entities but also deconstructs the query into a series of sub-problems. This multi-step decomposition facilitates improved LLM reasoning and consequently, enhanced performance.
@@
IV-B Comparison between human-centered path and turn-by-turn walking directions path
We compare human-centered guidance with conventional turn-by-turn directions routes (Figure 10). Turn-by-turn paths, generated from map topology, are constrained by predefined map data and may introduce unnecessary detours. In practice, pedestrians often prefer more direct or intuitive routes, such as open areas or informal shortcuts, reflecting a gap between algorithmic routing and human spatial reasoning. Human-centered guidance addresses this by pointing users toward the destination, allowing real-time adaptation based on their perception of walkable paths. As illustrated in Figure 11, our system integrates camera, geolocation, and orientation data to compute directional bearings and overlay AR cues. This method reduces detours, increases flexibility, and provides a more natural navigation experience in dynamic or unstructured environments where rigid step-by-step instructions may fail.
IV-C Evaluation on the embedding-based entity search
The XGBoost ranker was trained on a dataset of 500 image queries, each paired with manually annotated ground‑truth venues, and evaluated on a held‑out set of 50 queries. For benchmarking, we considered two baselines: (i) a distance‑based method that orders candidate places solely by geodesic proximity to the user, and (ii) a similarity‑based method that ranks candidates exclusively according to the cosine similarity between image and place embeddings. Table II summarizes the comparative performance of the proposed XGBoost ranker and two baseline methods, evaluated using Precision and Recall at Top‑1 and Top‑3. The Top-k Precision and Recall are defined as:
Top‑1 metrics () measure whether the highest‑ranked candidate is correct, providing a strict indicator of ranking accuracy at the very top. In contrast, Top‑3 metrics () evaluate the proportion and coverage of relevant items within the first three positions, reflecting the system’s ability to surface multiple correct candidates early in the ranking. As shown in Table II, the XGBoost ranker consistently outperforms both distance‑based and similarity‑based baselines across all metrics, with the largest gains observed in Top‑3 recall, indicating improved breadth of relevant retrieval without sacrificing top‑rank precision.
IV-D Evaluation on Spatial Understanding

### v4

@@ Abstract
Modern mapping tools remain fundamentally point–and–click, offering little support for interactive, multimodal engagement with either map views or real-world camera inputs. This gap becomes most apparent in the last-100-meter of navigation and during in-situ local discovery, where users naturally combine spatial reasoning over maps with ego-centric perception of their surroundings. We introduce IMAIA, an interactive geospatial assistance framework that unifies three core capabilities: (1) map-centric spatial understanding, (2) camera-to-place grounding, and (3) human-centered, bearing-aware navigation. IMAIA is composed of two interoperable components—Maps Plus and PAISA—coordinated by a lightweight multi-agent orchestrator. Maps Plus converts vector and satellite maps into a grid-aligned representation that enables flexible, view-conditioned spatial queries. PAISA fuses camera imagery with geospatial signals such as location, heading, and proximity to interpret ego-centric scenes and surface relevant attributes of nearby places. The framework is fully modular: vision-language backends are replaceable without changing system behavior. IMAIA provides responsive, interpretable, and context-aware assistance across both map-based and real-world interactions, enabling a new class of interactive mapping experiences that remain robust and practical for user-facing deployment.
1 Introduction
@@
Modern map applications remain largely point–and–click: users pan and zoom, then issue basic, limited, and inflexible queries. This interaction model breaks down for map-centric, view-conditioned queries—e.g., “What is the flower-shaped building next to the park in the top-right of what I am viewing?”—and for connecting what a mobile camera sees to the surrounding geospatial context. The desire to explore and navigate unfamiliar environments is fundamental, yet current tools remain inadequate: traditional travel and mapping systems, constrained by static methodologies, struggle with real-world dynamism—fluctuating conditions, imprecise signals, and unexpected disruptions that degrade the experience. Meanwhile, travel planning, navigation, and local discovery are typically engineered as isolated modules, yielding fragmented interactions and brittle handoffs [30, 10]; prior work even documents user behavior under disrupted plans [13], yet practical systems still fail to support coherent, interactive map understanding at the moment it is most needed.
@@
Recent advances in Large Language Models (LLMs) offer a path to transcend these limitations [16, 29, 1]. LLMs can process and synthesize diverse multimodal inputs, including text, imagery, geospatial data, and contextual cues [21, 12, 11], enabling more cohesive and adaptive geospatial assistants [28]. Two trends are particularly relevant. First, LLMs are increasingly adept at interpreting unstructured or ambiguous geospatial information, transforming vague user requests into precise map coordinates or actionable insights from noisy datasets [28]. Second, advancements in interactive and conversational AI highlight the need for multi-turn, context-aware interactions in navigation and discovery, where goals evolve over time [27]. In parallel, vision–language models (VLMs [31]) can describe images, but without explicit grounding to the current map state (viewport, scale, nearby entities) and geospatial signals (location, heading, proximity), their reasoning remains fragile in the presence of ambiguous visual cues or noisy inputs. [6, 8]. Prior work has shown that transferring knowledge from larger models to smaller task-specialized models through distillation can improve robustness and efficiency [7, 9], motivating our use of a distilled spatial reasoning module in IMAIA.
Against this backdrop, we focus on a concrete application problem: interactive, multimodal interaction with maps and real-world ego-centric views that bridges desktop trip planning and in-situ, last-100-meter navigation and local discovery. We target users in two complementary settings: (i) planning and exploring on a desktop or phone map, where they want to reason about what they see on vector and satellite views, and (ii) arriving near their destination, where they want to understand what their camera is facing in the real world and receive guidance that respects their current bearing and surroundings. To support this end-to-end experience, our system must unify three capabilities that are typically handled in isolation: (1) map-centric spatial understanding, where the map viewport is encoded as a quadkey-indexed grid and exposed as a structured visual prompt for interactive, view-conditioned reasoning over vector and satellite maps; (2) ego-centric scene understanding and geospatial place grounding, where first-person camera views are connected to nearby entities using both visual features and geospatial context—implemented through our PAISA module; and (3) human-centered navigation, providing bearing-aware guidance that matches how people naturally orient themselves and move through space in the last-100-meter phase.
In this paper, we introduce the Interactive Maps AI Assistant (IMAIA), a modular framework built around two tightly integrated components—Maps Plus and the Places AI Smart Assistant (PAISA)—coordinated by a lightweight multi-agent orchestration layer. At the system and application level, Maps Plus delivers an interactive map experience for desktop-style exploration over vector and satellite maps, while PAISA provides an AR-style experience for interacting with real-world scenes grounded in associated geospatial data. Concretely, IMAIA operationalizes the three capabilities above as follows:
Maps Plus: interactive map experience via quadkey-based visual prompting. We introduce Maps Plus, an interactive map interface that enables view-conditioned, map-centric queries directly over vector and satellite maps. The system converts the current viewport into a quadkey-indexed grid and attaches visual and semantic attributes to each tile, forming a structured visual prompt that supports deictic spatial reasoning during map exploration.
PAISA: ego-centric scene understanding and geospatial place grounding. We introduce PAISA, a camera-based module that grounds first-person views to nearby places by fusing visual input with geospatial signals such as location, heading, and proximity. Within a multi-agent architecture for orchestration, location intelligence, navigation, and spatial reasoning, PAISA connects observed scenes to surrounding entities and generates human-centered, bearing-aware navigation guidance.
IMAIA: a unified multimodal framework with strong empirical performance. To the best of our knowledge, IMAIA is the first framework that tightly couples interactive map exploration with camera-grounded place understanding and last-100-meter navigation within a unified multimodal AI assistant. The framework is model-agnostic and modular, enabling interchangeable VLM and vision backends. Empirically, Maps Plus improves place detection accuracy from under 43% to nearly 90%, while our spatial reasoning model achieves 84% accuracy with a 7.3 inference speedup over agent-based pipelines.
2 Related Work
2.1 Geospatial/Maps Intelligence with LLMs
@@
2.2 Spatial Intelligence and Reasoning with VLMs
@@
3 Interactive Maps AI Assistant
3.1 Maps Plus
Travelers usually begin by asking broad questions such as “What are the top locations to visit in X?”, “What should I see in X?”, or “Plan me a route through X.” Previous studies have shown that a large language model (LLM) can answer such queries either directly from its internal memory or by coordinating with external tools [28, 29], as illustrated in Figure 2. These approaches are referred to as Single Model, Model + Location, and Model + Verbose Location. Once users begin examining an interactive map, however, they often pose richer, map-centric questions that require spatial reasoning about what they are actually viewing rather than a static list of attractions. To meet these needs, we introduce a multimodal system that blends an LLM with image input and geospatial search (as shown in Figure 1). A user can click on map tiles or satellite imagery and converse naturally about that view, receiving answers grounded in both textual knowledge and visual context. As shown in Figure 2, our proposed method first determines the geographic focus and zoom level of the user’s current view, then scans the surrounding imagery on a regular grid to extract visual features and detect salient geographic entities. Finally, it queries a geospatial index with those entities and synthesizes the results so the LLM can craft an informed, location-aware response. By treating the map itself as conversational context, the system supports fluid trip planning and exploratory tasks that go well beyond what text-only approaches can deliver.
3.1.1 Location Awareness
@@
3.1.2 Quadkey-based Visual Prompting
Next, we convert the current map viewport into a quadkey-indexed [18] grid and overlay this grid on a simplified map, as illustrated in Figure 3. Each quadkey tile is treated as a structured element in a visual prompt [25] to the multimodal LLM: the model is tasked with identifying tiles that contain salient map entities such as roads, parks, or water bodies. This quadkey-based discretization associates detected entities with precise spatial locations on the map, enabling spatial correlation analysis and reasoning over tile-level relationships. By segmenting the map into these visually prompted tiles, the model can decompose complex spatial layouts into manageable units, making downstream tasks—such as answering questions about specific regions or patterns across the viewport—more accessible and interpretable.
3.1.3 Entity Search and Query Resolution
@@
3.2 Places AI Smart Assistant (PAISA)
@@
3.2.1 Interactive Navigation Agent
@@
3.2.2 Location Intelligence Agent
@@
3.2.3 Spatial Understanding Agent
Spatial reasoning is essential for connecting an ego-centric photo to nearby places during the final steps of navigation. Existing spatial VLMs (e.g., ASMv2 [23], SpatialVLM [2], SpatialRGPT [4]) provide strong relational understanding but are often too slow or misaligned with the fine-grained urban cues required in our application. To obtain an efficient and task-adapted module that integrates cleanly into the IMAIA framework, we distill GPT–4o into a Florence-2 model using a structured, multi-stage pipeline.
The agent takes a single street-level image and extracts (i) salient objects such as storefront signs, façades, or distinctive structural elements, and (ii) their spatial relationships, represented either as a scene graph or natural-language description. This representation supports two key system functions. First, when IMAIA retrieves cached street-level images of a destination, the agent generates relational descriptions (e.g., “the café entrance is just left of the red awning”), helping users visually confirm the location. Second, when users upload a photo, the agent converts it into a structured spatial record that downstream components use for grounding, matching, and contextual disambiguation. This pairing of destination-side and user-side spatial grounding improves interpretability and reliability in last-meter navigation.
Our distillation pipeline consists of three stages. In stage (i), GPT–4o-mini is prompted to extract candidate key entities for 40k street-view images, and the most frequent elements are retained as salient anchors. In stage (ii), YOLO-World [5] and Depth Anything V2 [26] provide 2D localization and depth cues, and GPT–4o generates pairwise spatial relations using Set-of-Mark-style prompting. In stage (iii), we build a supervised fine-tuning set by pairing each annotated image with diverse relational queries reflecting real urban navigation needs. We then fine-tune a Florence-2 model in its dense captioning format, yielding a compact and responsive spatial reasoning module suitable for real-time deployment within IMAIA.
3.2.4 Handle complex query with multi-agent reasoning
@@
4 Results
4.1 Place detection accuracy with Maps Plus
In this section, we evaluate and compare the quality of geospatial entity search across several methods. Specifically, we benchmark the Maps Plus approach against three widely used baselines (Figure 2) in Table 1: (1) Single Model: The model receives only the user query and the current map view image, relying solely on its pretrained knowledge. (2) Model + Location: The input includes the query, the map view image, and the geographic coordinates (latitude and longitude) of the map view. (3) Model with Verbose Location: Similar to the previous setup, but the location input is replaced with verbose descriptors such as city names and landmarks. In this study, we construct a dataset by selecting 10 cities across the United States and sampling points of interest (POIs) within a 20-kilometer radius of each city center. For each POI, we employ GPT-4o to generate synthetic queries based on information such as the POI’s attributes, geographic coordinates, and related contextual data. This process yields a total of 4,300 queries, for example, What is the lake at the top left part of the map.
@@ 4.1 Place detection accuracy with Maps Plus
Table 1 presents the accuracy achieved by various methods. The results indicate that our proposed approach attains significantly higher accuracy compared to the baselines, without requiring any fine-tuning of the LLM. The superior performance of the proposed method can be attributed to its efficient integration of grounding data, specifically entities retrieved from geospatial services, into the LLM’s processing. For instance, consider the query ”What is the coffee shop below the cinema?”. The proposed method not only furnishes a pertinent set of local entities but also deconstructs the query into a series of sub-problems. This multi-step decomposition facilitates improved LLM reasoning and consequently, enhanced performance.
@@
4.2 Comparison between human-centered path and turn-by-turn walking directions path
We compare human-centered guidance with conventional turn-by-turn (TBT) navigation (Figure 9). TBT routes rely on map topology and predefined road graphs, which may introduce unnecessary detours. In contrast, pedestrians often choose more direct paths, such as crossing open spaces or taking informal shortcuts, revealing a gap between algorithmic routing and human spatial reasoning. Human-centered guidance addresses this limitation by directing users toward the destination using their first-person camera view and orientation. As illustrated in Figure 10, our system integrates camera input, geolocation, and device orientation to compute destination bearings and overlay AR guidance cues. We evaluate this approach on 10 last-100-meter navigation scenarios. Four scenarios require a turn where the destination is not directly visible, and six involve destinations that are visible but may be temporarily occluded. Walking time is measured relative to standard TBT navigation. In turn-required scenarios, human-centered guidance reduces average walking time from 3.28 (min) to 2.08 (min) (63.5% of TBT). In directly visible scenarios, the reduction is larger, from 3.36 (min) to 1.07 (min) (32.1% of TBT). These results indicate that bearing-aware, scene-informed guidance can significantly reduce detours and improve navigation efficiency. The interface and interaction design were further refined through informal feedback from approximately 15 users, improving clarity and usability and resulting in the final interface presented in the paper.
4.3 Evaluation on the embedding-based entity search
The XGBoost ranker was trained on a dataset of 500 image queries, each paired with manually annotated ground‑truth venues, and evaluated on a held‑out set of 50 queries. For benchmarking, we considered two baselines: (i) a distance‑based method that orders candidate places solely by geodesic proximity to the user, and (ii) a similarity‑based method that ranks candidates exclusively according to the cosine similarity between image and place embeddings. Table 2 summarizes the comparative performance of the proposed XGBoost ranker and two baseline methods, evaluated using Precision and Recall at Top‑1 and Top‑3. The Top-k Precision and Recall are defined as:
Top‑1 metrics () measure whether the highest‑ranked candidate is correct, providing a strict indicator of ranking accuracy at the very top. In contrast, Top‑3 metrics () evaluate the proportion and coverage of relevant items within the first three positions, reflecting the system’s ability to surface multiple correct candidates early in the ranking. As shown in Table 2, the XGBoost ranker consistently outperforms both distance‑based and similarity‑based baselines across all metrics, with the largest gains observed in Top‑3 recall, indicating improved breadth of relevant retrieval without sacrificing top‑rank precision.
4.4 Evaluation on Spatial Understanding